# Optimal Multi-Distribution Learning

### v2

@@ 4.2 Measuring the trajectory of Hedge
As presented in Algorithm 2, there is a set of loss vectors $\mathcal{Y}$, and we want to solve the minimax problem
where the equivalence arises from von Neumann’s minimax theorem (v. Neumann,, 1928). By Hedge algorithm, we can obtain an $O(\varepsilon)$-minimax solution in $\frac{100\log(k)}{\varepsilon^{2}}$ rounds. The question is that:
  • Can we bound $\|\overline{w}^{T}\|_{1}=\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ by polylogarithmic terms?
To explain our two key techniques, we assume that
$\min_{y\in\mathsf{conv}(\mathcal{Y})}\max_{w\in\Delta(k)}w^{\top}y=\max_{w\in\Delta(k)}\min_{y\in\mathsf{conv}(\mathcal{Y})}w^{\top}y=0,$
Given this assumption, we then have $(w^{t})^{\top}y^{t}=\min_{y\in\mathcal{Y}}(w^{t})^{\top}y\leq 0,\forall t\in[T]$ and $\frac{1}{T}\sum_{t=1}^{T}(w^{t})^{\top}y^{t}\geq-C\epsilon$ for some constant $C$. In the following analysis, we assume $-(w^{t})^{\top}y^{t}=O(\epsilon)$ holds for any $t\in[T]$. While this assumption is not true in most cases, we can divide $[T]$ into $j_{\mathrm{max}}:=O\left(\log_{2}(1/\epsilon)\right)$ disjoint subsets as $\{\{t\in[T]:-(w^{t})^{\top}y^{t}\in(2^{-(j+1)},2^{j}]\}\}_{j=0}^{j_{\mathrm{max}}-2}$ and $\{t\in[T]:-(w^{t})^{\top}y^{t}\leq 2^{-(j_{\mathrm{max}}-1)}=O(\epsilon)\}$ and then work on each subset. As a trade-off, this operation leads to some additional logarithmic factors (see Lemma 29).
4.2.1 Doubling $w_{j}$ needs $\widetilde{\Omega}(1/\varepsilon^{2})$ steps
Instead of bounding $\|\overline{w}^{T}\|_{1}$ directly, our first attempt is to show that there are at most $\widetilde{O}(1)$ dimensions $i$ such that $\max_{1\leq t\leq T}w^{t}_{i}\geq c$, with $c>0$ some universal constant. For simplicity, we take $c=1/4$.
Let
$\mathcal{W}\coloneqq\Big{\{}i\in[k]\mid\max_{1\leq t\leq T}w_{i}^{t}\geq 1/4\big{\}}.$ (18)
Given that the stepsize $\eta$ is small enough, one can find, for each $i\in[k]$, an interval $[s_{i},e_{i}]$ such that
$\frac{1}{16}\leq w_{i}^{s_{i}}\leq\frac{1}{8},\qquad w_{i}^{e_{i}}\geq\frac{1}{4}\qquad\text{and}\qquad w_{i}^{t}\geq\frac{1}{8}$ (19)
for any $t$ obeying $s_{i}<t\leq e_{i}$. Then we would like to show that $e_{i}-s_{i}\geq\widetilde{\Omega}(1/\varepsilon^{2})$.
Since the stepsize $\eta=O(\varepsilon)$, we can easily get that $e_{i}-s_{i}=\widetilde{\Omega}(1/\varepsilon)$. Meanwhile, when running the Hedge algorithm for adversarial bandits, the adversarial could choose the loss properly so that $e_{i}-s_{i}=\widetilde{\Theta}(1/\varepsilon)$. Fortunately, we can focus on a bilinear game where the opponent plays best response. By definition, for any $1\leq t\leq\tau\leq T$,
$(w^{t})^{\top}y^{t}\leq 0,\quad(w^{t})^{\top}y^{\tau}\geq(w^{t})^{\top}y^{t}.$
For any $1\leq t_{1}\leq t_{2}\leq T$, we then have
$(w^{t_{1}})^{\top}\sum_{\tau=t_{1}}^{t_{2}-1}y^{\tau}\geq(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}.$ (20)
By observing that $\sum_{\tau=t_{1}}^{t_{2}-1}y_{i}^{\tau}=\frac{1}{\eta}\log\left(\frac{W^{t_{2}}_{i}}{W^{t_{1}}_{i}}\right)$, we have
$\sum_{i=1}^{k}w_{i}^{t_{1}}\log\left(\frac{W_{i}^{t_{2}}}{W_{i}^{t_{1}}}\right)\geq\eta(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}.$ (21)
On the other hand, since the optimal minimax value is 00, we can show that
$\log\left(\frac{\sum_{i=1}^{k}W_{i}^{t_{2}}}{\sum_{i=1}^{k}W_{i}^{t_{1}}}\right)=\sum_{\tau=t_{1}}^{t_{2}-1}\log\left(\frac{\sum_{i=1}^{k}W_{i}^{\tau+1}}{\sum_{i=1}^{k}W_{i}^{\tau}}\right)\leq O\big{(}(t_{2}-t_{1})\eta^{2}\big{)}.$ (22)
Putting (21) and (22) together, we arrive at
$-\sum_{i=1}^{k}w_{i}^{t_{1}}\log\left(\frac{w_{i}^{t_{2}}}{w_{i}^{t_{1}}}\right)\leq O\left(\eta^{2}(t_{2}-t_{1})-\eta(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}\right)=O(\eta^{2}(t_{2}-t_{1})),$ (23)
which is exactly
$\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}\leq O\big{(}\eta^{2}(t_{2}-t_{1})\big{)}.$ (24)
Finally, we obtain that
$t_{2}-t_{1}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}}{\varepsilon^{2}}\right).$ (25)
By setting $t_{1}=s_{i}$ and $t_{2}=e_{i}$, we can deduce that
$e_{i}-s_{i}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}}{\varepsilon^{2}}\right)=\Omega\left(\frac{1}{\varepsilon^{2}}\right).$ (26)
Also noting that for any $t$, there are at most $8$ dimensions $i$ such that $s_{i}\leq t\leq e_{i}$, it holds that
@@ 4.2.1 Doubling $w_{j}$ needs $\widetilde{\Omega}(1/\varepsilon^{2})$ steps
$8T\geq\sum_{i\in\mathcal{W}}(e_{i}-s_{i})\geq|\mathcal{W}|\cdot\Omega\left(\frac{1}{\varepsilon^{2}}\right),$ (27)
which implies that $|\mathcal{W}|$ is at most $O(T\varepsilon^{2})=O(\log(k))$.
In the analysis above, we analyze the property of the trajectory of Hedge algorithm with exact loss vector. However, in Algorithm 1, we need to sample to estimate the loss vector. In this case, (24) would be replaced by
$\mathsf{KL}(w^{t_{1}}\,\|\,w^{t_{2}})\leq O(\eta^{2}(t_{2}-t_{1}))+O\left(\eta\sum_{\tau=t_{1}}^{t_{2}-1}(w^{t_{1}}+w^{\tau})^{\top}\Delta^{\tau}\right),$ (28)
where $\Delta^{\tau}:=\left(\widehat{y}^{\tau}-y^{\tau}\right)$ with $\widehat{y}^{\tau}$ as the stochastic estimation of $y^{\tau}$. A naive bound for the second term in right hand side of (28) would be $\widetilde{O}\left(\eta\sqrt{(t_{2}-t_{1})}\right)$ by Azuma’s inequality. However, this bound is not sufficient because when $(t_{2}-t_{1})\varepsilon^{2}\ll 1$, the term $\widetilde{O}\left(\eta\sqrt{(t_{2}-t_{1})}\right)$would be the leading term. To address this problem, we consider weighted sampling to construct the empirical loss vector. As presented in Line 1 Algorithm 1, we manage to reduce the variance of $\Delta^{\tau}$ properly so that $\left(\eta\sum_{\tau=t_{1}}^{t_{2}-1}(w^{t_{1}}+w^{\tau})^{\top}\Delta^{\tau}\right)$ is roughly bounded by $O(\eta^{2}(t_{2}-t_{1}))$.
4.2.2 Dealing with the segments
In Section 4.2.1, we show that for any $1\leq t_{1}\leq t_{2}\leq T$
$t_{2}-t_{1}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}}{\varepsilon^{2}}\right).$ (29)
By this inequality, we bound the size of $\mathcal{W}$ by $O(\log(k))$. A natural idea is to generalize the arguments to bound the size of
$\mathcal{W}(p):=\left\{i\in[k]|\max_{1\leq t\leq T}w_{i}^{t}\in[2p,4p]\right\}$
for $p\in[0,1]$. However, the arguments above can not provide a desired bound for $\mathcal{W}(p)$ when $p$ is small. For each $i\in\mathcal{W}(p)$, let $[s_{i},e_{i}]$ be the interval such that $\frac{p}{2}\leq w_{i}^{s_{i}}\leq p$, $w_{i}^{e_{i}}\geq 2p$ and $w_{i}^{t}\geq p$ for any $s_{i}<t\leq e_{i}$. By (29), we learn that
$e_{i}-s_{i}\geq\Omega\left(\frac{\mathsf{KL}\big{(}\mathrm{Ber}(p)\,\|\,\mathrm{Ber}(2p)\big{)}}{\varepsilon^{2}}\right)\geq\Omega\left(\frac{p}{\varepsilon^{2}}\right),$ (30)
where $\mathrm{Ber}(q)$ denote the Bernoulli distribution with mean $q$. Following the arguments in (27), we hope to bound
$\sum_{i\in\mathcal{W}(p)}(e_{i}-s_{i})=O(T).$ (31)
However, this might be not true because for each $t\in[T]$, there are possibly more than $\frac{1}{2p}$ coordinates $i\in\mathcal{W}(p)$ such that $t\in[s_{i},e_{i}]$, where we can at most derive $|\mathcal{W}(p)|\leq O(T\varepsilon^{2}/p^{2})$. As a result, the contribution of coordinates in $\mathcal{W}(p)$ to the target $\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ is $\widetilde{O}(1/p)$, which could be much larger than $\widetilde{O}(1)$ when $p$ is small.
To address this problem, we note that if there are some coordinates $i\in\mathcal{W}(p)$ sharing similar $[s_{i},e_{i}]$, we can obtain a better bound. For example, suppose that for any $i\in\overline{\mathcal{W}}(p)\subset\mathcal{W}(p)$, $s_{i}=s$ and $e_{i}=e$, we then have that
@@ 4.2.2 Dealing with the segments
$e-s\geq\Omega\left(\frac{|\overline{\mathcal{W}}(p)|p}{\varepsilon^{2}}\right),$ (32)
which is much better compared to the original bound $\Omega\left(\frac{p}{\varepsilon^{2}}\right)$ if $|\overline{\mathcal{W}}(p)|$ is large.
By this observation, we consider to merge the dimensions in $\mathcal{W}(p)$ with similar $[s_{i},e_{i}]$. To facilitate analysis, we introduce the notation of segment.
@@
For $p,x>0$ and $i\in[k]$, we say $(t_{1},t_{2})$ is a $(p,q,x)$-segment iff there exists $\mathcal{I}\subset[K]$ such that $\sum_{i\in\mathcal{I}}w^{t_{1}}_{i}\in\left[\frac{p}{2},p\right]$, $\sum_{i\in\mathcal{I}}w^{t_{2}}_{i}\geq p\exp(x)$ and $\sum_{i\in\mathcal{I}}w_{i}^{t}\geq q$ for any $t_{1}\leq t\leq t_{2}$. We name $t_{1}$ as the starting point and $t_{2}$ as the ending point. And we also call $\mathcal{I}$ as the index set.
By choosing $p/4\leq q\leq p/2$, a $(p,q,x)$-segment captures the variation of probabilities from $p$ to $p\exp(x)$. We say two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$ are disjoint iff $s_{1}<e_{1}\leq s_{2}<e_{2}$ or $s_{2}<e_{2}\leq s_{1}<e_{1}$. Using (29), it is easy to derive $t_{2}-t_{1}\geq\Omega(px^{2}/\varepsilon^{2})$ for a $(p,p/2,x)$-segment $(t_{1},t_{2})$.
For each $i\in\mathcal{W}(p)$, there exists a $\left(\frac{p}{4},\frac{p}{8},\log(2)\right)$-segment $(s_{i},e_{i})$ with index set =$\mathcal{I}=\{i\}$. It then follows that $e_{i}-s_{i}\geq\Omega\left(\frac{p}{\varepsilon^{2}}\right)$ for each $i\in\mathcal{W}(p)$. As a result, if the segments $(s_{i},e_{i})$ are mutually disjoint, $|\mathcal{W}(p)|$ is at most $O(\log(k)/p)$. More generally, if we can divide the $|\mathcal{W}(p)|$ segments into disjoint blocks such that the segments in the same block share the same starting and ending points, we can also derive that $|\mathcal{W}(p)|\leq O(\log(k)/p)$. Taking Figure 1 as an example. Suppose we have $\ell$ blocks and the $i$-th block (with starting-ending points as $(\widetilde{s}^{i},\widetilde{e}^{i})$) contains $m_{i}$ coordinates satisfying that $\tilde{s}^{1}<\tilde{e}^{1}\leq\tilde{s}^{2}<\tilde{e}^{2}\leq\dots\leq\tilde{s}^{\ell}<\tilde{e}^{\ell}$. Then $(\widetilde{s}^{i},\widetilde{e}^{i})$ constructs a $\left(m_{i}\cdot\frac{p}{4},m_{i}\cdot\frac{p}{8},\log(2)\right)$-segment with index set as the $m_{i}$ coordinates in the $i$-th block, which means that
$\widetilde{e}^{i}-\widetilde{s}^{i}\geq\Omega\left(\frac{m_{i}p}{\varepsilon^{2}}\right).$
Taking sum over $1\leq i\leq\ell$, we obtain
$|\mathcal{W}(p)|\leq\sum_{i=1}^{\ell}m_{i}\leq O\left(\sum_{i=1}^{\ell}\frac{(\widetilde{e}^{i}-\widetilde{s}^{i})\varepsilon^{2}}{p}\right)\leq O\left(\frac{T\varepsilon^{2}}{p}\right)=O\left(\frac{\log(k)}{p}\right),$ (33)
which implies that $\sum_{i\in\mathcal{W}(p)}\max_{1\leq t\leq T}w_{i}^{t}\leq O(\log(k))$. With standard doubling arguments, it follows that $\sum_{i\in[k]}\max_{1\leq t\leq T}w_{i}^{t}=O(\log^{2}(k))$.
By this observation, we introduce the concept of a regular configuration (see an example in Figure 1).
Definition 3 (Configuration).
A configuration $\mathsf{Conf}$ is a set of intervals $\mathsf{Conf}=\{[a_{i},b_{i}]\}_{i=1}^{m}$ with $b_{i}>a_{i}$ for each $i\in[m]$ (repeatable elements are allowed). We say the configuration $\mathsf{Conf}$ is regular iff for any $i,j\in[m]$, one of the three following claims holds.
  • $a_{i}<b_{i}\leq a_{j}<b_{j}$;
  • $a_{j}<b_{j}\leq a_{i}<b_{i}$;
  • $a_{i}=a_{j}$, $b_{i}=b_{j}$.
In words, (33) states that if $\{[s_{i},e_{i}]\}_{i\in\mathcal{W}(p)}$ forms a regular configuration, we then have $|\mathcal{W}(p)|=O\left(\frac{\log(k)}{p}\right)$ However, the general configuration of the segments might be irregular because it is possible that two segments are intersecting (see Figure 2). That is, for two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$, it holds that $s_{1}<s_{2}<e_{1}<e_{2}$. In this case, the approach above can at most bound $e_{2}-s_{1}$ by $\max\{e_{2}-s_{2},e_{1}-s_{1}\}$ instead of the desired bound $(e_{1}-s_{1})+(e_{2}-s_{2})$. To address this problem, we consider to construct a regular configuration with sub-segments [^5] A sub-segment is equivalent to a sub-interval of a segment. We refer to Figure 6 for more details. of the original segments.
Our first step is to align one side of the segments. As presented in Figure 3, we divide the whole learning process into disjoint blocks such that the segments in each block have a common inner point. In the meantime, a segment is discarded if it intersects with more than one blocks. We show that there exists a regular configuration of the blocks such that at most $(1-\frac{1}{3(\log_{2}(T)+1)})\cdot|\mathcal{W}(p)|$ segments are discarded. In other words, at least $\frac{1}{3(\log_{2}(T)+1)}\cdot|\mathcal{W}(p)|$ segments are contained by these blocks. Since the blocks are disjoint to each other, it suffices to operate on one block. Without loss of generality, we assume that there is only one block, which also means that all segments have a common inner point $t_{\mathrm{mid}}$. Then we divide this block into two parts according to $t_{\mathrm{mid}}$ (see Figure 4). Since
$\log\left(\frac{w_{i}^{e_{i}}}{w_{i}^{t_{\mathrm{mid}}}}\right)+\log\left(\frac{w_{i}^{t_{\mathrm{mid}}}}{w_{i}^{s_{i}}}\right)\geq\log(2),$
it then holds either $(s_{i},t_{\mathrm{mid}})$ is a $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment with index ${i}$ or $(t_{\mathrm{mid}},e_{i})$ is a $\left(w_{i}^{t_{\mathrm{mid}}},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment with index set $\{i\}$. Since $w_{i}^{t_{\mathrm{mid}}}\in[p/4,p/2]$, we can roughly view $(t_{\mathrm{mid}},e_{i})$ as a $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment in the latter case. Therefore, we can find at least $\frac{|\mathcal{W}(p)|}{3(\log_{2}(T)+1)}\cdot\frac{1}{2}$ different $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$ segments with the same starting (ending) points.
Without loss of generality, we assume that these $\frac{|\mathcal{W}(p)|}{6(\log_{2}(T)+1)}\cdot$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segments share a common starting point. As presented in Figure 5, we denote the common starting point as $e_{0}$, and re-order the dimensions so that $e_{i}$ is non-decreasing in $i$. Then we construct $O(\log_{2}(k))$ regular configurations using a recursion (see Lemma 28) , and show that at least one of these configurations contains $\widetilde{O}\left(|\mathcal{W}(p)|\right)$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segments.
In Figure 6, we present a toy example in the case $k=8$. In this toy example We in total have $5$ configurations with 5 different colors. For each $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment in Figure 5, it contributes an $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segment to at least one of these configurations. By the pigeonhole princeton, there must be one regular configuration with $\frac{|\mathcal{W}(p)|}{6(\log_{2}(T)+1)(\log_{2}(k)+2)}$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segments (see Lemma 28).
Figure 1: Regular configuration.
Figure 2: General irregular configuration.
Figure 3: Partition of blocks.
Figure 4: Aligning one side of the configuration. The unfilled part of the segments mean that the variation over $w_{i}^{s_{i}}\to w_{i}^{t_{\mathrm{mid}}}$ (i.e., $\log\left(w_{i}^{t_{\mathrm{mid}}}/w_{i}^{s_{i}}\right)$) is not significant enough.
Figure 5: A group of segments with common starting point.
Figure 6: Construction of the regular configurations. Each segment is cut into at most $\log_{2}(k)+1$ sub-segments with different colors.
5 Analysis (proof of Theorem 1)
Theorem 1 follows by combining Lemma 4, Lemma 5 and Lemma 6.
Recall that $\varepsilon_{1}=\frac{\varepsilon}{100}$. In Lemma 4, we show that, with high-probability,
$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (34)
holds for any $1\leq t\leq T$. Assuming (34) holds for any $1\leq t\leq T$, we continue by proving that $h^{\mathsf{final}}$ is $\varepsilon$-optimal in Lemma 5, and providing a bound for the sample complexity in Lemma 6.

### v3

@@ 4.1 Towards sample reuse: uniform concentration and a key quantity
This algorithm is often referred to as the Hedge algorithm, which is known to yield an $\varepsilon$-minimax solution within $O\big{(}\frac{\log(k)}{\varepsilon^{2}}\big{)}$ iterations. A challenging question relevant to our analysis is:
  • Question: can we bound $\|\overline{w}^{T}\|_{1}\coloneqq\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ in Algorithm 2 by poly-logarithmic terms?
As it turns out, we can answer this question affirmatively (see Lemma 3), and the key ideas will be elucidated in the remainder of this section.
4.2.1 First attempt: bounding the number of distributions with $\max_{1\leq t\leq T}w^{t}_{i}=\Omega(1)$
Instead of bounding $\|\overline{w}^{T}\|_{1}$ directly, our first attempt is to look at those $i\in[k]$ with large $\max_{1\leq t\leq T}w^{t}_{i}$ (more specifically, $\max_{1\leq t\leq T}w^{t}_{i}\geq 1/4$) and show that:
  • there exist at most $\widetilde{O}(1)$ coordinates $i\in[k]$ obeying $\max_{1\leq t\leq T}w^{t}_{i}\geq 1/4$ (or some other universal constant).
In other words, we would like to demonstrate that the cardinality of the following set is small:
$\mathcal{W}_{\mathsf{large}}\coloneqq\big{\{}i\in[k]\mid\max\nolimits_{1\leq t\leq T}w_{i}^{t}\geq 1/4\big{\}}.$ (18)
To do so, note that some standard “continuity”-type argument tells us that: for a sufficiently small stepsize $\eta$, one can find, for each $i\in\mathcal{W}_{\mathsf{large}}$, a time interval $[s_{i},e_{i}]\subseteq[0,T]$ obeying
$1/16\leq w_{i}^{s_{i}}\leq 1/8,\qquad w_{i}^{e_{i}}\geq 1/4,\qquad\text{and}\qquad w_{i}^{t}\geq 1/8~{}~{}\forall t\in(s_{i},e_{i}].$ (19)
In words, $w_{i}^{t}$ at least doubles from $t=s_{i}$ to $t=e_{i}$. We claim for the moment that
$e_{i}-s_{i}\geq{\Omega}(1/\eta^{2})={\Omega}(1/\varepsilon^{2})\qquad\forall i\in\mathcal{W}_{\mathsf{large}}.$ (20)
Additionally, observe that for any $t$, there exist at most $8$ coordinates $i\in\mathcal{W}_{\mathsf{large}}$ such that $s_{i}\leq t\leq e_{i}$ (since $w_{i}^{t}\geq 1/8$ for every $t\in[s_{i},e_{i}]$). This reveals that
@@ 4.2.1 First attempt: bounding the number of distributions with $\max_{1\leq t\leq T}w^{t}_{i}=\Omega(1)$
$8T\geq\sum_{i\in\mathcal{W}_{\mathsf{large}}}(e_{i}-s_{i})\geq|\mathcal{W}_{\mathsf{large}}|\cdot\Omega\left(1/\varepsilon^{2}\right),$ (21)
which combined with our choice of $T=\widetilde{O}(1/\varepsilon^{2})$ (cf. line 1 of Algorithm 1) yields
$|\mathcal{W}_{\mathsf{large}}|\leq O(T\varepsilon^{2})=\widetilde{O}(1).$
In words, the number of distributions with large $\max_{1\leq t\leq T}w^{t}_{i}$ (i.e., $\max_{1\leq t\leq T}w^{t}_{i}\geq 1/4$) is fairly small.
Proof sketch for (20).
Let us briefly discuss the high-level proof ideas. Following standard analysis for the Hedge algorithm (e.g., shalev2012online; lattimore2020bandit), we can often obtain (under certain mild conditions)
$\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}\leq O\big{(}\eta^{2}(e_{i}-s_{i})\big{)}.$
Combine this relation with basic properties about the KL divergence (see, e.g., Lemmas 10 and 11 in Appendix A) and the choice $\eta=\varepsilon/10$ to obtain
$e_{i}-s_{i}$ $\geq\Omega\big{(}\eta^{-2}\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}\big{)}\geq\Omega\Big{(}\eta^{-2}\mathsf{KL}\big{(}\mathsf{Ber}(w_{i}^{s_{i}})\,\|\,\mathsf{Ber}(w_{i}^{e_{i}})\big{)}\Big{)}$
$\geq\Omega\bigg{(}\eta^{-2}\frac{\frac{1}{16}\cdot 2^{2}}{4}\bigg{)}=\Omega\left(1/\eta^{2}\right)=\Omega\left(1/\varepsilon^{2}\right).$
4.2.2 More general cases: issues and solutions
Naturally, one would hope to generalize the arguments in Section 4.2.1 to cope with more general cases. More specifically, let us look at the following set
$\mathcal{W}(p)\coloneqq\left\{i\in[k]\mid\max\nolimits_{1\leq t\leq T}w_{i}^{t}\in[2p,4p]\right\},$ (22)
defined for each $p\in[0,1]$. If one could show that
$|\mathcal{W}(p)|=\widetilde{O}(1/p)\qquad\text{for each }p,$ (23)
then a standard doubling argument would immediately lead to
$\|\overline{w}^{T}\|_{1}$ $=\sum_{i\in[k]}\max_{1\leq t\leq T}w_{i}^{t}\approx\sum_{j=1}^{\log_{2}k}\sum_{i\in\mathcal{W}(2^{-j})}\max_{1\leq t\leq T}w_{i}^{t}\leq\sum_{j=1}^{\log_{2}k}4\cdot 2^{-j}\cdot\big{|}\mathcal{W}(2^{-j})\big{|}$
$\leq\sum_{j=1}^{\log_{2}k}4\cdot 2^{-j}\cdot\widetilde{O}\bigg{(}\frac{1}{2^{-j}}\bigg{)}=\widetilde{O}(1).$
A technical issue.
Nevertheless, simply repeating the arguments in Section 4.2.1 fails to deliver the desirable bound (23) on $|\mathcal{W}(p)|$ when $p$ is small. Briefly speaking, for each $i\in\mathcal{W}(p)$, let $[s_{i},e_{i}]$ represent a time interval (akin to (19)) such that
$p/2\leq w_{i}^{s_{i}}\leq p,\qquad w_{i}^{e_{i}}\geq 2p\qquad\text{and}\qquad w_{i}^{t}\geq p\quad\text{for any }s_{i}<t\leq e_{i}.$ (24)
Repeating the heuristic arguments in Section 4.2.1 leads to
$e_{i}-s_{i}\geq\Omega\big{(}\eta^{-2}\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}\big{)}\geq\Omega\Big{(}\eta^{-2}\mathsf{KL}\big{(}\mathsf{Ber}(w_{i}^{s_{i}})\,\|\,\mathsf{Ber}(w_{i}^{e_{i}})\big{)}\Big{)}\geq\Omega(\eta^{-2}p)=\Omega\left(p/\varepsilon^{2}\right).$ (25)
Given that each $t$ is contained within at most $1/(p/2)$ intervals associated with $\mathcal{W}(p)$, repeat the arguments for (21) to derive
@@ 4.2.2 More general cases: issues and solutions
$\frac{1}{p/2}\cdot T\geq\sum_{i\in\mathcal{W}(p)}(e_{i}-s_{i})\geq|\mathcal{W}(p)|\cdot\Omega\left(p/\varepsilon^{2}\right)\qquad\Longrightarrow\qquad|\mathcal{W}(p)|\leq\widetilde{O}\bigg{(}\frac{1}{p^{2}}\bigg{)}.$
This bound, however, is clearly loose compared to the desirable one in (23).
Our solution.
To address this issue, we make two key observations below that inspire our approach:
  • Shared intervals. For the interval $[s_{i},e_{i}]$ associated with $i$ as defined above, if there exist other indices sharing the same interval $[s_{i},e_{i}]$ (in the sense that relations analogous to (24) are satisfied for other indices), then it is plausible to improve the bound. For instance, if a set $\mathcal{M}_{i}\subseteq[k]$ of indices share the same interval $[s_{i},e_{i}]$, then one can follow the heuristic argument in (25) to obtain
    $e_{i}-s_{i}\geq\Omega\big{(}\eta^{-2}\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}\big{)}\geq\Omega\left(\eta^{-2}\mathsf{KL}\bigg{(}\mathsf{Ber}\Big{(}\sum_{j\in\mathcal{M}_{i}}w_{j}^{s_{i}}\Big{)}\,\|\,\mathsf{Ber}\Big{(}\sum_{j\in\mathcal{M}_{i}}w_{j}^{e_{i}}\Big{)}\bigg{)}\right)\geq\Omega\left(p|\mathcal{M}_{i}|/\varepsilon^{2}\right),$ (26)
    which clearly strengthens the original bound (25) if $|\mathcal{M}_{i}|$ is large.
  • Disjoint intervals. Consider the special case where $\mathcal{W}(p)$ can be divided into subsets $\{\mathcal{V}_{n}\}_{n=1}^{N}$ obeying
    • (i)
      for each $n\in[N]$, all indices in $\mathcal{V}_{n}$ share the same interval $[s_{n},e_{n}]$ (defined analogously as (24));
    • (ii)
      the intervals $\{[s_{n},e_{n}]\}_{n=1}^{N}$ are disjoint.
    Then one can derive the desired bound on $|\mathcal{W}(p)|$. More precisely, it follows from (26) that
    $e_{n}-s_{n}\geq\Omega\left(p|\mathcal{V}_{n}|/\varepsilon^{2}\right),$ (27)
    which together with the disjointness property yields
    $|\mathcal{W}_{j}|=\sum_{n=1}^{N}|\mathcal{V}_{n}|\leq\sum_{n=1}^{N}O\bigg{(}\frac{(e_{n}-s_{n})\varepsilon^{2}}{p}\bigg{)}\leq O\bigg{(}\frac{T\varepsilon^{2}}{p}\bigg{)}=\widetilde{O}\bigg{(}\frac{1}{p}\bigg{)}.$ (28)
In light of the above discussion, it is helpful to (a) merge those indices that share similar intervals, and (b) identify disjoint intervals whose associated indices can cover a good fraction of $\mathcal{W}_{j}$.
Motivated by the aforementioned observation about “shared intervals,” we introduce the notion of “segments” to facilitate analysis.
@@
For any $p,x>0$ and $i\in[k]$, we say that $(t_{1},t_{2})$ is a $(p,q,x)$-segment if there exists a subset $\mathcal{I}\subseteq[k]$ such that
  • $\mathrm{(i)}$
    $\sum_{i\in\mathcal{I}}w^{t_{1}}_{i}\in\left[p/2,p\right]$,
  • $\mathrm{(ii)}$
    $\sum_{i\in\mathcal{I}}w^{t_{2}}_{i}\geq p\exp(x)$,
  • $\mathrm{(iii)}$
    $\sum_{i\in\mathcal{I}}w_{i}^{t}\geq q$ for any $t_{1}\leq t\leq t_{2}$.
We shall refer to $t_{1}$ as the starting point and $t_{2}$ as the end point, and call $\mathcal{I}$ the associated index set. Moreover, two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$ are said to be disjoint if $s_{1}<e_{1}\leq s_{2}<e_{2}$ or $s_{2}<e_{2}\leq s_{1}<e_{1}$.
This definition allows one to pool indices with similar intervals together.
In general, however, it is common for two segments to be overlapping (see Figure 2 in Appendix C.5), which precludes us from directly invoking our aforementioned observation about “disjoint intervals.” To address this issue, our strategy is to extract out shared sub-segments [^7] A sub-segment refers to a sub-interval of a segment, as illustrated in Figure 6. of (a subset of) these segments in a meticulous manner. Encouragingly, it is possible to find such sub-segments that taken collectively cover a good fraction (i.e., $\frac{1}{\mathrm{poly}\log(k,T)}$) of all segments, meaning that we do not have to discard too many segments. The construction is built upon careful analysis of these segments, and will be elucidated in Appendix C.
5 Analysis for VC classes (proof of Theorem 1)
The main steps for establishing Theorem 1 lie in proving three key lemmas, as stated below.
The first lemma is concerned with the hypothesis $h^{t}=\arg\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})$ (cf. line 1 of Algorithm 1); in words, $h^{t}$ is the minimizer of the empirical loss function $\widehat{L}^{t}(\cdot,w^{t})$, computed using samples obtained up to the $t$-th round. The following lemma tells us that: even though $h^{t}$ is an empirical minimizer, it almost optimizes the weighted population loss $L(\cdot,w^{t})$. In other words, this lemma justifies that the adaptive sampling scheme proposed in Algorithm 1 ensures faithfulness of the empirical loss and its minimizer; here, we recall that $\varepsilon_{1}=\varepsilon/100$ (cf. line 1 of Algorithm 1).